eviant:

PLS HELP!!! Which is the equation of a line that is perpendicular to the line represented by y=3/4 x -1/2?

eviant:

@Allison @BRAINIAC

BRAINIAC:

Keep in mind that, when both eqn of lines are perpendicular to each other,the slope is not the same and u would need to use \(m_1m_2=-1\) to find the other slope of eqn of line

BRAINIAC:

\(m_1=\frac{3}{4}\) sub the value into the formula to find \(m_2\)

Allison:

-4/3x?

BRAINIAC:

\(m_1m_2=-1\) \(\frac{3}{4}m_2=-1\) \(m_2=\frac{-4}{3}\)

Allison:

:DD

BRAINIAC:

Therefore,the eqn of line will be \(y=-\frac{4}{3}x\)

Allison:

YAY

BRAINIAC:

:) ;)

Allison:

(: